OpenStudy (anonymous):

what volume of 0.3682 M H2SO4 solution is required to react with 0.4198g of Al(CN)3 according to the reaction, 2Al(CN)3 + 3H2SO4 --> 6HCN + Al2(SO4)3? the answer is 16.28mL, but I don't know how to solve this, please help!!

OpenStudy (aaronq):

Have you done other stoichiometry problems before?

OpenStudy (anonymous):

no this is the first time, i am taking it for summer

OpenStudy (aaronq):

hm okay. We'll the goal is to first convert your reactants into moles. For this problem, you need to know the molar mass of \(Al(CN)_3\). Then, to find the moles, use: \(moles=\dfrac{mass}{molar~mass}\) I'll wait till you do that to explain the rest.

OpenStudy (anonymous):

ok I got 105.01 for the molar mass

OpenStudy (aaronq):

okay, now find the moles.

OpenStudy (anonymous):

so I wrote down 0.004 moles, 0.4198g/105.01g/mol

OpenStudy (aaronq):

okay. now, we want to know how many moles of \(H_2SO_4\) we need. To do so, we need a ratio of the moles of each reactant and their stoichiometric coefficients (from the balanced reaction). n is the symbol for moles \(\dfrac{n_{H2SO_4}}{3}=\dfrac{n_{Al(CN_3)}}{2}\) Plug in your values (moles of aluminium cyanide) and solve for moles of sulfuric acid: \(\dfrac{n_{H2SO_4}}{3}=\dfrac{0.004}{2}\rightarrow n_{H2SO_4}=\dfrac{3*0.004}{2}=0.006 ~moles\) good so far?

OpenStudy (anonymous):

yes

OpenStudy (aaronq):

okay, so lastly, they want the answer in terms of volume, given a concentration in Molarity. So we use the Molarity (M) formula: \(M=\dfrac{n_{solute}}{L_{solution}}\)

OpenStudy (aaronq):

Solve for volume (which is in liters here)

OpenStudy (anonymous):

yay i got the answer!! thank you so much, i finally understood it! :D

OpenStudy (aaronq):

no problem! glad i could help (just remember that in stoichiometry you need to convert what you're given to moles, then work from there)

OpenStudy (anonymous):

yes sir!